OpenStudy (anonymous):

A 70.0 L sample of gas at 20.0ºC and 600. mm Hg expands to 90.0 L at 15.0ºC. What is the new gas pressure?

OpenStudy (whpalmer4):

\[PV=nRT\]to the rescue again, although we can simplify it to \[\frac{PV}{T} = k\] Remember to convert the temperature into degrees Kelvin! Then we can restate the problem as \[\frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2}\]plug in the values and find the value of \(P_2\)